OpenStudy (anonymous):

How this solution possible ? \(\lim_{n \to \infty}\frac{n^{2}+n-2}{n^{2}+n} = \lim_{n \to \infty}(1-\frac{2}{n^{2}+n})\)

OpenStudy (anonymous):

I cant understand how we get 1 ?

OpenStudy (agent47):

(n^2+n-2)=(n-1)(n+2)

OpenStudy (agent47):

n^2+n=n(n+1)

OpenStudy (agent47):

Ohh wait

OpenStudy (anonymous):

ok..

OpenStudy (agent47):